When 25.0 mL of 0.500 M H2SO4 is added to 25.0 mL of 1.00 M KOH in a coffee-cup calorimeter at
23.50
1
Expert's answer
2021-10-11T06:12:32-0400

v=0.5/25=0.02mol

v=1/25=0.04mol

v=0.04+0.02=0.06mol

C=23.5*0.06=1.41mol/ml